OpenStudy (anonymous):

Solve -3x2 + 30x - 90 = 0.

OpenStudy (amoodarya):

\[ax^2+bx+c=0 \\x_1,x_2=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] in this case \[b^2-4ac<0\]

OpenStudy (anonymous):

thank u